# The nucleon axial radius

### v2

@@ Abstract
We present the first systematic study of the relativistic axial-vector four-current distributions inside a nucleon. We show in particular that the slope of the axial form factor $G_{A}(Q^{2})$ in the forward limit – conventionally denoted as $R^{2}_{A}$ in the literature – does not represent the three-dimensional mean-square axial radius in the Breit frame, but corresponds instead to a contribution to the mean-square spin radius. We derive explicit expressions for the latter in different frames and find in general additional contributions that depend on both the nucleon mass and the forward values of the axial-vector form factors $G_{A}(0)$ and $G_{P}(0)$. This provides an additional key motivation for on-going lattice QCD calculations and future experimental measurements of the induced pseudoscalar form factor $G_{P}(Q^{2})$.
Introduction. A fundamental aspect of the nucleon spin structure is captured by the axial-vector form factors (FFs), which are Lorentz-invariant quantities playing a crucial role in the weak sector, especially in neutrino scattering Mann et al. (1973); Baker et al. (1981); Horstkotte et al. (1982); Ahrens et al. (1987); Aguilar-Arevalo et al. (2010, 2015); Sufian et al. (2020); Cai et al. (2023) and oscillation Ayres et al. (2007); Abe et al. (2011); Abi et al. ; Abe et al. ; An et al. (2016) experiments. Numerous theoretical investigations of these FFs based on chiral perturbation theory and phenomenological models can be found in Refs. Bernard et al. (1994); Schindler et al. (2007); Eichmann and Fischer (2012); Dahiya and Randhawa (2014); Anikin et al. (2016); Mamedov et al. (2017); Mondal et al. (2020); Chen et al. (2022); Cheng et al. (2022); Liu et al. (2023); Borah et al. (2024). Tremendous progress from the lattice QCD side has also been reported in the last few years Green et al. (2017); Shintani et al. (2019); Jang et al. (2020); Bali et al. (2020); Park et al. (2022); Alexandrou et al. (2021); Djukanovic et al. (2022); Jang et al. (2024); Alexandrou et al. (2024). For recent reviews on axial-vector FFs, see e.g. Refs. Bernard et al. (2002); Gorringe and Fearing (2004); Meyer et al. (2022); Sajjad Athar et al. (2023); Gupta (2024).
@@
in agreement with older analyses Yennie et al. (1957); Friar and Negele (1975) and the requirement that the total electric charge must transform as a Lorentz scalar Lorcé (2020). The BF energy being given by $P^{0}_{B}=p^{\prime 0}_{B}=p^{0}_{B}=M\sqrt{1+\tau}$ with $\tau\equiv Q^{2}/(4M^{2})$, one concludes that the relativistic mean-square charge radius of the proton is in fact given by Chen and Lorcé (2023)
@@
The additional term, known as the Darwin-Foldy term, actually appeared in Ref. Ernst et al. (1960) but was later discarded by hand by the authors to comply with the conventions in force in atomic spectroscopy.
@@
In Figure 1, we illustrate the spin distribution (12) inside the nucleon using the corresponding axial-vector FFs $G_{A}(Q^{2})\equiv\sum_{f}G_{A}^{f}(Q^{2})$ and $G_{P}(Q^{2})\equiv\sum_{f}G_{P}^{f}(Q^{2})$ from recent lattice QCD calculations Alexandrou et al. (2021), where $G_{X}^{f}(Q^{2})$ for $X=A,P$ and $f=u,d$ are parametrized as
@@
with the dipolar parameters given in Table 1. It should be noted that the contributions from heavier quarks (e.g., $s$ and $c$ quarks) are not included in our numerical calculations due to their small contributions but large uncertainties Alexandrou et al. (2021). We find in particular that $r_{S}\approx 0.916~{}\text{fm}$ and $R_{A}\approx 0.567~{}\text{fm}$ by using the lattice data from Ref. Alexandrou et al. (2021). This indicates that the relativistic contribution $\frac{1}{4M^{2}}\left(1+\frac{2\,G_{P}(0)}{G_{A}(0)}\right)$ in Eq. (13) is not small.
@@
The 2D mean-square helicity (or axial) radius is then naturally defined on the LF for $\lambda^{\prime}=\lambda$ as Diehl and Hagler (2005)
@@
Since LF distributions are two-dimensional, it is tempting to interpret $R^{2}_{A}$ as the corresponding three-dimensional mean-square axial radius, e.g. based on the Abel-Radon transformation Panteleeva and Polyakov (2021); Kim and Kim (2021). This would be justified if we could demonstrate that the LF axial charge distribution is the projection of a spherically symmetric 3D distribution onto the transverse plane. Unfortunately, 3D LF distributions with strict probabilistic interpretation do not exist because of the necessary $\Delta^{+}=0$ constraint Freese and Miller (2022).
@@
Alternatively, one can consider the longitudinal LF spin distribution given by $S_{L,\text{LF}}\equiv\frac{1}{4}\text{Tr}[\sigma_{z}\,\mathcal{A}^{+}_{\text{LF}}]$, which is nothing other than half of the light-front helicity distribution. Since by definition the mean-square radius does not depend on the overall normalization of the distribution, we obtain naturally
@@
If we use instead the transverse LF spin distribution $S_{T,\text{LF}}\equiv\frac{1}{4}\text{Tr}[\boldsymbol{\sigma}_{\perp}\cdot\boldsymbol{\mathcal{A}}_{\perp,\text{LF}}]=\hat{\boldsymbol{s}}_{\perp}\cdot\boldsymbol{S}_{T,\text{LF}}$, where the numerical result for $\boldsymbol{S}_{T,\text{LF}}(\boldsymbol{b}_{\perp};P^{+})$ is illustrated in Fig. 2, we find
@@
Note that the $P^{+}$-dependence of $S_{T,\text{LF}}$ enters merely as a global factor and drops out when calculating the mean-square radius. The difference between longitudinal and transverse LF mean-square spin radii is of relativistic origin, and disappears in the static limit $M\to\infty$.
@@
When $P_{z}>0$ and the target is longitudinally polarized, the axial density is nonzero. We can then define a meaningful 2D mean-square axial radius in a generic EF for a longitudinally polarized target $s^{\prime}=s$:
@@
Alternatively, we can consider the longitudinal and transverse EF spin distributions $S_{L,\text{EF}}\equiv\frac{1}{4}\text{Tr}[\sigma_{z}\,\mathcal{A}_{z,\text{EF}}]$ and $S_{T,\text{EF}}\equiv\frac{1}{4}\text{Tr}[\boldsymbol{\sigma}_{\perp}\cdot\boldsymbol{\mathcal{A}}_{\perp,\text{EF}}]=\hat{\boldsymbol{s}}_{\perp}\cdot\boldsymbol{S}_{T,\text{EF}}$, where the numerical result for $\boldsymbol{S}_{T,\text{EF}}(\boldsymbol{b}_{\perp};P_{z})$ is illustrated in Fig. 3. The corresponding 2D mean-square radii read
@@
In the infinite-momentum limit $P_{z}\to\infty$, we recover, as expected, the LF mean-square spin radii (19) and (20). Interestingly, if we average over all spin directions and take $P_{z}\to 0$, we find
@@
Conclusions. We investigated in detail in this Letter whether the widespread interpretation of $R_{A}^{2}$ given by Eq. (1) as the nucleon mean-square axial radius is justified or not.
In the Breit frame, where 3D spatial distributions are traditionally defined, the axial density vanishes identically (11). This means that the 3D mean-square nucleon axial radius does not exist, and thus the interpretation of $R_{A}$ as the nucleon axial radius does not rest on sound physical grounds. In contrast, we demonstrated that a physically meaningful 3D nucleon mean-square spin radius (13) can be unambiguously defined in terms of the same axial-vector FFs.
Another option is to consider 2D spatial distributions in the light-front formalism. For a longitudinally polarized nucleon, the 2D mean-square axial radius is equal to the mean-square longitudinal spin radius, and is given by $\frac{2}{3}R^{2}_{A}$ (19). For a transversely polarized nucleon, however, the mean-square axial radius vanishes while the mean-square transverse spin radius receives an additional contribution on top of the $\frac{2}{3}R^{2}_{A}$ term (20).
An interpolation between the Breit frame and light-front results can be obtained by using the concept of 2D elastic frame distributions, based on the phase-space formalism. When the nucleon has finite average momentum $P_{z}$, the 2D mean-square axial and transverse spin radii receive additional $P_{z}$-dependent contributions compared to the light-front ones, see Eqs. (26) and (27). These extra contributions formally disappear, however, in the infinite-momentum limit $P_{z}\to\infty$.
In summary, contrary to what is usually stated in the literature we did not find any proper justification for interpreting the quantity $R^{2}_{A}$ as the physical 3D mean-square axial radius of a spin-$\frac{1}{2}$ target (a nucleon for instance). On the other hand, the nucleon mean-square spin radius $\langle r_{\text{spin}}^{2}\rangle$ (13) is a well-defined and physically meaningful 3D mean-square radius that characterizes more realistically the spatial extension of the weak content of the nucleon. From a 2D perspective, however, the quantity $\frac{2}{3}R^{2}_{A}$ can consistently be interpreted as the 2D mean-square axial or longitudinal spin radius, albeit only for a longitudinally polarized spin-$\frac{1}{2}$ target.
While our analysis does not directly impact the phenomenology of axial-vector form factors, it explicitly demonstrates that the quantity $R^{2}_{A}$ alone does not faithfully measure the spatial extension of the weak content of the nucleon. According to Eq. (13), the 3D mean-square spin radius $\langle r_{\text{spin}}^{2}\rangle$ requires besides $R_{A}^{2}$ the additional determination of the ratio $G_{P}(0)/G_{A}(0)$, and provides a key motivation for further lattice QCD calculations and experimental measurements of the induced pseudoscalar form factor $G_{P}(Q^{2})$ in (anti)neutrino-nucleon scatterings.

### v3

# Nucleon axial radius

@@ Abstract
We present the first systematic study of the relativistic axial-vector four-current distributions inside a nucleon. We show in particular that the slope of the axial form factor $G_{A}(Q^{2})$ in the forward limit—conventionally denoted as $R^{2}_{A}$ in the literature—does not represent the three-dimensional mean-square axial radius in the Breit frame, but corresponds instead to a contribution to the mean-square spin radius. We derive explicit expressions for the latter in different frames and find in general additional contributions that depend on both the nucleon mass and the forward values of the axial-vector form factors $G_{A}(0)$ and $G_{P}(0)$. This provides an additional key motivation for ongoing lattice quantum chromodynamics (QCD) calculations and future experimental measurements of the induced pseudoscalar form factor $G_{P}(Q^{2})$.
Introduction. A fundamental aspect of the nucleon spin structure is captured by the axial-vector form factors (FFs), which are Lorentz-invariant quantities playing a crucial role in the weak sector, especially in neutrino scattering Mann et al. (1973); Baker et al. (1981); Horstkotte et al. (1982); Ahrens et al. (1987); Aguilar-Arevalo et al. (2010a, b); Fields et al. (2013); Aguilar-Arevalo et al. (2015); Meyer et al. (2016); Sufian et al. (2020); Cai et al. (2023) and oscillation Ayres et al. (2007); Abe et al. (2011); Abi et al. ; Abe et al. ; An et al. (2016) experiments. Numerous theoretical investigations of these FFs based on chiral perturbation theory and phenomenological models can be found in Refs. Bernard et al. (1994); Schindler et al. (2007); Eichmann and Fischer (2012); Dahiya and Randhawa (2014); Anikin et al. (2016); Mamedov et al. (2017); Mondal et al. (2020); Chen et al. (2022); Cheng et al. (2022); Liu et al. (2023); Borah et al. (2024); Ramalho et al. . Tremendous progress from the lattice quantum chromodynamics (QCD) side has also been reported in the last few years Green et al. (2017); Shintani et al. (2019); Jang et al. (2020); Bali et al. (2020); Park et al. (2022); Alexandrou et al. (2021); Djukanovic et al. (2022); Jang et al. (2024); Alexandrou et al. (2024). For recent reviews on axial-vector FFs, see e.g. Refs. Bernard et al. (2002); Gorringe and Fearing (2004); Meyer et al. (2022); Sajjad Athar et al. (2023); Gupta (2024).
@@
in agreement with older analyses Yennie et al. (1957); Friar and Negele (1975) and the requirement that the total electric charge must transform as a Lorentz scalar Lorcé (2020). The BF energy being given by $P^{0}_{B}=p^{\prime 0}_{B}=p^{0}_{B}=M\sqrt{1+\tau}$ with $\tau\equiv Q^{2}/(4M^{2})$, one concludes that the relativistic mean-square charge radius of the proton is in fact given by Chen and Lorcé (2023); Friar et al. (1997)
@@
The additional term, known as the Darwin-Foldy term Yennie et al. (1957); Foldy and Wouthuysen (1950); Foldy (1952); Friar et al. (1997), see also the Eqs. (5b) and (5c) of Ref. Friar et al. (1997), actually appeared in Ref. Ernst et al. (1960) but was later discarded by hand by the authors to comply with the conventions in force in atomic spectroscopy Miller (2019).
@@
In Fig. 1, we illustrate the 3D spin distribution (12) inside the nucleon in the $r_{y}=0$ plane using the nucleon axial-vector FFs $G_{A}(Q^{2})\equiv\sum_{f}G_{A}^{f}(Q^{2})$ and $G_{P}(Q^{2})\equiv\sum_{f}G_{P}^{f}(Q^{2})$ from recent lattice QCD calculations Alexandrou et al. (2021), where $G_{X}^{f}(Q^{2})$ for $X=A,P$ and $f=u,d$ are parametrized as
@@
with the dipolar parameters given in Table 1. It should be noted that the contributions from heavier quarks (e.g., $s$ and $c$ quarks) are not included in our numerical calculations due to their small contributions but large uncertainties Alexandrou et al. (2021). We find in particular that $r_{S}\approx 0.916~{}\text{fm}$ and $R_{A}\approx 0.567~{}\text{fm}$ by using the lattice parametrization of $G_{X}^{f}(Q^{2})$ from Ref. Alexandrou et al. (2021). This indicates that the relativistic contribution $\frac{1}{4M^{2}}\left(1+\frac{2\,G_{P}(0)}{G_{A}(0)}\right)$ in Eq. (13) is not small.
@@
The 2D transverse mean-square helicity (or axial) radius is naturally defined on the LF for $\lambda^{\prime}=\lambda$ as Diehl and Hagler (2005)
@@
Since LF distributions are two-dimensional, it is tempting to interpret $R^{2}_{A}$ as the corresponding 3D mean-square axial radius, e.g. based on the Abel-Radon transformation Panteleeva and Polyakov (2021); Kim and Kim (2021). This would be justified if we could demonstrate that the LF axial charge distribution is the projection of a spherically symmetric 3D distribution onto the transverse plane. Unfortunately, 3D LF distributions with strict probabilistic interpretation do not exist because of the necessary $\Delta^{+}=0$ constraint Freese and Miller (2022).
@@
Alternatively, one can consider the longitudinal LF spin distribution given by $S_{L,\text{LF}}\equiv\frac{1}{4}\text{Tr}[\sigma_{z}\,\mathcal{A}^{+}_{\text{LF}}]$, which is nothing other than half of the light-front helicity distribution. Since by definition the (transverse) mean-square radius does not depend on the overall normalization of the distribution, we obtain naturally
@@
If we use instead the transverse LF spin distribution $S_{T,\text{LF}}\equiv\frac{1}{4}\text{Tr}[\boldsymbol{\sigma}_{\perp}\cdot\boldsymbol{\mathcal{A}}_{\perp,\text{LF}}]=\hat{\boldsymbol{s}}_{\perp}\cdot\boldsymbol{S}_{T,\text{LF}}$, where the numerical result of $\boldsymbol{S}_{T,\text{LF}}(\boldsymbol{b}_{\perp};P^{+})$ is illustrated in Fig. 2, we find
@@
Note that the $P^{+}$-dependence of $S_{T,\text{LF}}$ enters merely as a global factor and drops out when calculating the transverse mean-square radius. The difference between longitudinal and transverse LF (transverse) mean-square spin radii is of relativistic origin (20), and disappears in the static limit $M\to\infty$.
@@
When $P_{z}>0$ and the target is longitudinally polarized, the axial density is nonzero. We can then define a meaningful 2D transverse mean-square axial radius in a generic EF for a longitudinally polarized target $s^{\prime}=s$:
@@
Alternatively, we can consider the longitudinal and transverse EF spin distributions $S_{L,\text{EF}}\equiv\frac{1}{4}\text{Tr}[\sigma_{z}\,\mathcal{A}_{z,\text{EF}}]$ and $S_{T,\text{EF}}\equiv\frac{1}{4}\text{Tr}[\boldsymbol{\sigma}_{\perp}\cdot\boldsymbol{\mathcal{A}}_{\perp,\text{EF}}]=\hat{\boldsymbol{s}}_{\perp}\cdot\boldsymbol{S}_{T,\text{EF}}$, where the numerical result of $\boldsymbol{S}_{T,\text{EF}}(\boldsymbol{b}_{\perp};P_{z})$ is illustrated in Fig. 3. The corresponding 2D transverse mean-square radii read
@@
In the infinite-momentum limit $P_{z}\to\infty$, we recover, as expected, the LF transverse mean-square spin radii (19) and (20). Interestingly, if we average over all spin directions and take $P_{z}\to 0$, we find
@@
Conclusions. We investigated in detail in this letter whether the widespread interpretation of $R_{A}^{2}$ given by Eq. (1) as the nucleon mean-square axial radius is justified or not. In the Breit frame, where 3D spatial distributions are traditionally defined, the axial density vanishes identically (11). This means that the 3D mean-square nucleon axial radius does not exist, and thus the interpretation of $R_{A}$ as the nucleon axial radius does not rest on sound physical grounds. In contrast, we demonstrated that a physically meaningful 3D nucleon mean-square spin radius (13) can be unambiguously defined in terms of the same axial-vector FFs.
Another option is to consider 2D spatial distributions in the light-front formalism. For a longitudinally polarized nucleon, the 2D transverse mean-square axial radius is equal to the transverse mean-square longitudinal spin radius, and is given by $\frac{2}{3}R^{2}_{A}$ (19). For a transversely polarized nucleon, however, the transverse mean-square axial radius vanishes while the transverse mean-square transverse spin radius receives an additional contribution on top of the $\frac{2}{3}R^{2}_{A}$ term (20).
An interpolation between the Breit frame and light-front results can be obtained by using the concept of 2D elastic frame distributions, based on the phase-space formalism. When the nucleon has finite average momentum $P_{z}$, the 2D transverse mean-square axial and transverse spin radii receive additional $P_{z}$-dependent contributions compared to the light-front ones, see Eqs. (26) and (27). These extra contributions formally disappear, however, in the infinite-momentum limit $P_{z}\to\infty$.
In summary, contrary to what is usually stated in the literature we did not find any proper justification for interpreting the quantity $R^{2}_{A}$ as the physical 3D mean-square axial radius of a spin-$\frac{1}{2}$ target (a nucleon for instance). On the other hand, the 3D mean-square spin radius $\langle r_{\text{spin}}^{2}\rangle$ (13) is a well-defined and physically meaningful radius that characterizes more realistically the spatial extension of the weak content of the system. From a 2D perspective, however, the quantity $\frac{2}{3}R^{2}_{A}$ can consistently be interpreted as the 2D transverse mean-square axial or longitudinal spin radius, albeit only for a longitudinally polarized spin-$\frac{1}{2}$ target.
While our analysis does not directly impact the phenomenology of axial-vector form factors, it explicitly demonstrates that the quantity $R^{2}_{A}$ alone does not faithfully measure the spatial extension of the weak content of the nucleon. According to Eq. (13), the 3D mean-square spin radius $\langle r_{\text{spin}}^{2}\rangle$ requires besides $R_{A}^{2}$ an additional determination of the ratio $G_{P}(0)/G_{A}(0)$, which provides an additional key motivation for ongoing lattice QCD calculations and future experimental measurements in (anti)neutrino-nucleon scatterings of the induced pseudoscalar form factor $G_{P}(Q^{2})$.